Light bulbs are measured in lumens (light output), watts (energy used), and hours (life). A soft white light bulb has a mean life of 700 hours and a standard deviation of 35 hours. Julia claims that the soft light bulbs she uses give a mean life of 800 hours. Use the z-score approach to test wether her claim is true or not.
Answer by stanbon(75887)

Light bulbs are measured in lumens (light output), watts (energy used), and hours (life).
A soft white light bulb has a mean life of 700 hours and a standard deviation of 35 hours.
Julia claims that the soft light bulbs she uses give a mean life of 800 hours. Use the z-score approach to test whether her claim is true or not.
-----
Find the probability the life is >= 800 if u = 700 and sigma = 35
z(800) = (800-700)/35 = 2.857
----
P(life >=800) = P(z > 2.857) = 0.0021...
-----
Chances are 2 out of a thousand that Julia is correct.
============================================
Cheers,
Stan H.